OpenStudy (anonymous):

Solve system of equation with matrix. X1 - X2 - X3 = 1 -2X1 +6X2 + 10X3 = 14 2X1 +X2 +6X3

OpenStudy (e.mccormick):

Know how to put it in matrix form?

OpenStudy (anonymous):

Yes, is there a special format for on here?

OpenStudy (e.mccormick):

You can do it in \(\LaTeX\) if you know how. I can do that too.

OpenStudy (anonymous):

1 -1 -1 1 -2 6 10 14 2 1 6 9

OpenStudy (anonymous):

ok?

OpenStudy (e.mccormick):

\(\left[ \begin{array}{ccc|c} 1 & -1 & -1 & 1\\ -2 & 6 & 10 & 14\\ 2 & 1 & 6 & 9\\ \end{array}\right ]\)

OpenStudy (e.mccormick):

OK, know your elementary row operations?

OpenStudy (anonymous):

Just learning it......can I start with R2 = R2 + R3?

OpenStudy (e.mccormick):

That is a good start.

OpenStudy (e.mccormick):

\(\left[ \begin{array}{ccc|c} 1 & -1 & -1 & 1\\ 0 & 7 & 16 & 23\\ 2 & 1 & 6 & 9\\ \end{array}\right ]\)

OpenStudy (e.mccormick):

If you want to know what I typed, it is: ``` \left[ \begin{array}{ccc|c} 1 & -1 & -1 & 1\\ 0 & 7 & 16 & 23\\ 2 & 1 & 6 & 9\\ \end{array}\right ] ``` Makes it all pretty.

OpenStudy (e.mccormick):

Well, that between `\(` and `\)`.

OpenStudy (anonymous):

ok..

OpenStudy (e.mccormick):

So, what do you think the next step would be?

OpenStudy (anonymous):

(-2)R1 + R3 = R3

OpenStudy (e.mccormick):

And what would you get?

OpenStudy (anonymous):

1 -1 -1 1 0 7 16 23 0 3 8 7

OpenStudy (e.mccormick):

Yep. And now it gets fun.

OpenStudy (e.mccormick):

There are two ways to go from that point. You can use a LCM type solution to eliminate more, ir you can work with fractions.

OpenStudy (anonymous):

Im suppose to use Guass Jordan.......Im not sure what you mean.

OpenStudy (e.mccormick):

This is Guass Jordan. I am talking about the algebra needed to eliminate thing in the second row.

OpenStudy (anonymous):

Im not really sure where to go from here...

OpenStudy (e.mccormick):

I meant second column. Your first cokumn is done. You have a sole row with a number in it. This is sometimes called a pivot. Anyhow, you have 1 alone in the first row, first column. Now, you need to get the second row, second column to 1 and the other rows in the second column to 0.

OpenStudy (anonymous):

R2 = R2 - 2R3 ?

OpenStudy (e.mccormick):

Oh, that works nice. Avoids the fractions.

OpenStudy (anonymous):

so... 1 -1 -1 1 0 1 0 9 0 3 8 7

OpenStudy (e.mccormick):

OK. They set that up nice. So now use your new R2 to eliminate the 3 in R3C2

OpenStudy (anonymous):

that would give.... 1 -1 -1 1 0 1 0 9 0 0 8 -20

OpenStudy (e.mccormick):

Yep... and you need to do the same sort of thing to R1. Get rid of C2 there.

OpenStudy (anonymous):

can I use R1 + R2 = R1 ?

OpenStudy (e.mccormick):

Yep.

OpenStudy (anonymous):

that would give... 1 0 -1 10 0 1 0 9 0 0 0 2

OpenStudy (anonymous):

sorry thats wrong..

OpenStudy (e.mccormick):

Yah. LOL

OpenStudy (anonymous):

1 0 -1 10 0 1 0 9 0 0 2 -5

OpenStudy (e.mccormick):

OK. Now you just need to work with the last row. No way to avoid the fractions now.

OpenStudy (anonymous):

OK. But after that how do I take care of the -1 R1 C3?

OpenStudy (anonymous):

lol sorry I see it.

OpenStudy (anonymous):

It seems that these are always fairly tedious problems? Or something for a program to think about for large matrices?

OpenStudy (e.mccormick):

Well, there are some tools you learn later that can help, and there is software to do the work. However, matrices allow you to do things that would be exceptionally hard without them. Like Google. It is a really huge matrix. Or rotating pictures. Or encryption.

OpenStudy (anonymous):

Google is a matrix?

OpenStudy (e.mccormick):

Yes. They use a form of hastable lookup database, and that is based on matrixes.

OpenStudy (e.mccormick):

So, after you divide R3 by 2, and that gets you R3, and then add it to R1, what is your final matrix?

OpenStudy (anonymous):

1 0 0 25/2 0 1 0 9 0 0 1 5/2

OpenStudy (e.mccormick):

25/2?

OpenStudy (anonymous):

15/2

OpenStudy (e.mccormick):

Oh, I see. You forgot it was -5 on the last line. You also missed it in the answer. So yah: \(\left[ \begin{array}{ccc|c} 1 & 0 & 0 & \dfrac{15}{2}\\ 0 & 1 & 0 & 9\\ 0 & 0 & 1 & - \dfrac{5}{2}\\ \end{array}\right ]\) \(x_1 + 0 + 0 = \dfrac{15}{2}\) \(0 + x_2 + 0 = 9\) \(0 + 0 + x_3 = -\dfrac{5}{2}\) So there you are.

OpenStudy (e.mccormick):

And that is Gauss-Jordan.

OpenStudy (e.mccormick):

Here is another one I worked for someone that was interested in learning how to do these. It has a link to a PDF on the methods: http://openstudy.com/users/e.mccormick#/updates/5275c700e4b00c1700b26a41

OpenStudy (e.mccormick):

And if you end up doing any Matrix Multiplication, this was my small writeup on that: http://assets.openstudy.com/updates/attachments/515c4ceee4b0507ceba37899-e.mccormick-1369021610530-matrixmultiplication.png

OpenStudy (anonymous):

Thank you!